OpenStudy (anonymous):

Simplify. -(-11) - |-12| + (-6) =

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

-(-11) = 11 - I12I = -12 + (-6) = -6 -------------- so 11-12-6 = -7

OpenStudy (anonymous):

5-|-12|